OpenStudy (mollynicole2015):

Solve for f. -6f + 9 = -3f

pooja195 (pooja195):

Add 6t to both sides

pooja195 (pooja195):

-3+6=?

OpenStudy (mollynicole2015):

3

pooja195 (pooja195):

Goot both just divide both sides by 3t 9/3=3t/3 t=?

OpenStudy (mollynicole2015):

3

OpenStudy (mollynicole2015):

Man you're quick. :D

pooja195 (pooja195):

:)

OpenStudy (mollynicole2015):

Thanks Pooj. <3

pooja195 (pooja195):

np :)

OpenStudy (vanilla_ice03):

wooooop!...... sorry lolz

OpenStudy (mollynicole2015):

XD